two friends leave a hotel at the same time traveling in opposite direction. They travel for four hours and are then 480 miles apart. If Lily travels 10 miles per our faster than Noe, find the average rate of speed for each person.
Answer by mananth(16946)

Noe speed = x mph
lily speed = x+10 mph
they are moving away from each other
add up speeds
x+x+10
=2x+10
time = 4 hours
distance = 480 miles
d=rt
480 = 4(2x+10)
/4
120=2x+10
120-10=2x
110=2x
/2
x=55
55 & 65 are the speeds of Noe & lily repectively